# A Hassle-free Algorithm for Private Learning in Practice: Don’t Use Tree Aggregation, Use BLTs

### v1

@@ Abstract
The state-of-the-art for training on-device language models for mobile keyboard applications combines federated learning (FL) with differential privacy (DP) via the DP-Follow-the-Regularized-Leader (DP-FTRL) algorithm. Two variants of DP-FTRL are used in practice, tree aggregation and matrix factorization. However, tree aggregation suffers from significantly suboptimal privacy/utility tradeoffs, while matrix mechanisms require expensive optimization parameterized by hard-to-estimate-in-advance constants, and high runtime memory costs. This paper extends the recently introduced Buffered Linear Toeplitz (BLT) mechanism to multi-participation scenarios. Our BLT-DP-FTRL maintains the ease-of-use advantages of TreeAgg, while essentially matching matrix factorization in terms of utility and privacy. We evaluate BLT-DP-FTRL on the StackOverflow dataset, serving as a re-producible simulation benchmark, and across four on-device language model tasks in a production FL system. Our empirical results highlight the advantages of the BLT mechanism and elevate the practicality and effectiveness of DP in real-world scenarios.
@@ 2.3 Matrix Factorization (MF) for DP-FTRL
$\widetilde{{\bm{A}}{\bm{X}}}={\bm{B}}({\bm{C}}{\bm{X}}+\zeta{\bm{Z}})={\bm{A}}{\bm{X}}+\zeta{\bm{C}}^{-1}{\bm{Z}}\Rightarrow\tilde{\Delta}^{t}=\Delta^{t}+(\zeta{\bm{C}}^{-1}{\bm{Z}})_{t,:}$ (1)
@@ 2.4 BLT Mechanisms in DP-FTRL
Applying the parameterization in Eq. 2 (also appear in [14, Sec 5]) to [14, Alg 1], and initializing buffers ${\bm{S}}_{-1}\leftarrow\mathbf{0}\in\mathbb{R}^{d\times m}$, we can efficiently compute ${\bm{Z}}_{t,:}$ from $\hat{\bm{Z}}_{t,:}$ in the streaming setting, ${\bm{Z}}_{t,:}=\hat{\bm{Z}}_{t,:}+\omega^{T}{\bm{S}}_{t-1},{\bm{S}}_{t}=\operatorname{diag}(\theta){\bm{S}}_{t-1}+\bm{1}_{d}\hat{\bm{Z}}_{t,:}$. We rearrange the update equations to get $\hat{\bm{Z}}$ from ${\bm{Z}}$ and ${\bm{S}}$,
@@
For completeness, we provide the streaming multiplication algorithm of ${\bm{Z}}={\bm{C}}\hat{\bm{Z}}$ and $\hat{\bm{Z}}=C^{-1}{\bm{Z}}$ for ${\bm{C}}=\text{BLT}(\theta,\omega)$ in Alg. 2 and Alg. 3, respectively. Alg. 2 is a direct application of [14, Alg 1] and only used to derive Alg. 3, which is our streaming algorithm for generating correlated noise with BLTs using only $dm$ memory. Finally, we apply the streaming correlated noise $\hat{\bm{Z}}$ by BLT from Alg. 3 (using Eq. 3) in Alg. 1 for the BLT-DP-FTRL algorithm, i.e.,
@@
       $\theta,\omega$ for ${\bm{C}}=\text{BLT}(\theta,\omega)$
@@
  Initialize buffers ${\bm{S}}_{-1}\leftarrow\mathbf{0}\in\mathbb{R}^{m\times d}$
@@
       $\theta,\omega$ for ${\bm{C}}=\text{BLT}(\theta,\omega)$
@@
  Initialize buffers ${\bm{S}}_{-1}\leftarrow\mathbf{0}\in\mathbb{R}^{m\times d}$
@@ 3.1 Sensitivity Under Multiple Participations
The sensitivity of general Toeplitz matrices with decaying coefficients is recently discussed in Kalinin and Lampert [24, Thm. 2], which we restate in Thm. 3.1 with our notation. The participation pattern $\pi^{\star}$ simply puts the $k$ participations as early as possible, with each participation separated by exactly $b$. This sensitivity computation is important for both DP accounting and optimizing for BLT parameters in Sec. 3.3.
@@ 3.2 Evaluating Matrix Factorization Mechanisms
Eqs. 9 and 10 measure the distribution of noise to approximate the privacy-utility trade-off as the loss, which do not depend on the noise multiplier $\alpha=\sigma/\operatorname{sens}_{\Pi}({\bm{C}})$ that is directly used in accounting for the DP guarantees, as discussed in [14, Introduction]. For optimized ${\bm{C}}$ in matrix factorization mechanisms (e.g., TreeAgg, BandMF, and BLT), specific DP guarantees are achieved by scaling $\alpha$ (and corresponding $\sigma$ in Alg. 1). The total noise on the prefix sum $BZ$ also scales MaxLoss and RmsLoss by $\alpha$. Hence, without loss of generality, we use $RmsLoss$ and $MaxLoss$ to compare the optimality of different matrix factorization mechanisms, which is equivalent to assuming $\alpha=1$ that corresponds to $(\epsilon=5.3,\delta=10^{-7})$-DP (for example).
Note we deviate from the definitions of [14, arXiv v3], in order to distinguish the error introduced by ${\bm{B}}$ from the total loss (which is what we optimize and use to compare mechanisms), which incorporates the sensitivity of ${\bm{C}}$.
@@ 3.3 Optimizing BLTs for Multiple Participations
  $p(x)=\prod_{i\in[d]}(1-\theta_{i}x)$
  $q(x)=\prod_{i\in[d]}(1-\hat{\theta}_{i}x)$
@@
For the typical scale of $n<10^{5}$ in FL systems, rather than deriving closed forms for sensitivity and error as in Dvijotham et al. [14], we use an alternative approach that is flexible and simple to implement. Recall the properties of BLTs discussed in Sec. 2.4, we parameterize the optimization of the BLT by the pair $(\theta,\hat{\theta})$. Dvijotham et al. [14, Lem. 5.2] implies given a pair $(\theta,\hat{\theta})$, there exist unique $(\omega,\hat{\omega})$ such the $\text{BLT}(\theta,\omega)^{-1}=\text{BLT}(\hat{\theta},\hat{\omega})$, and we can compute $\omega$ and $\hat{\omega}$ in time $\mathcal{O}(d^{2})$; this result is summarized below as Alg. 5. Thus, given a $(\hat{\theta},\theta)$, we can efficiently compute the Toeplitz coefficients of ${\bm{C}}$ (using Eq. 2 applied to $(\theta,\omega)$) and ${\bm{C}}^{-1}$ (applying Eq. 2 to $(\hat{\theta},\hat{\omega})$). From the Toeplitz coefficients of ${\bm{C}}$ we can then efficiently compute sensitivity using Thm. 3.1. RmsError can be computed efficiently from the Toepltiz coefficients of ${\bm{C}}^{-1}$ following the approach of McKenna [25, Prop. 3.1], and a simple generalization of this approach applies to to MaxError as well. For completeness we summarize in the following proposition:
@@ 5 Production LMs for Mobile Keyboard
We summarize the privacy and utility results for es-ES, id-ID, pt-BR, and pt-PT LMs in Tab. 4, and show the privacy-utility curves for training pt-PT in Fig. 7. We provide additional curves for other LMs in Figs. 13, 14 and 15 inApp. B, and discuss the observations here. (1) Most of the models achieve DP guarantee of $\epsilon<10$ with exception of $\epsilon~{}10$ for pt-PT due to the challenge of small population; the pt-BR model trained with BLT-16.1 achieves $\epsilon<1$ at round 2000.DP guarantees of $\epsilon<10$ is commonly used for machine learning, and $\epsilon<1$ are considered strong guarantees [29]. To achieve single-digit DP guarantees in practice without sacrificing the utility, the production LMs are trained with large number of clients per round (known as report goal in FL systems). We use typical report goal 6.5K [35] for es-ES, id-ID and pt-BR, and use a smaller report goal 3K for pt-PT with a smaller population. We additionally run BLT-DP-FTRL with larger report goal and linearly scale up the noise multiplier to keep the signal-to-noise ratio for utility: BLT-16.1 with report goal 12K for pt-BR and BLT-5.5 with report goal 6K for pt-PT. The resulting min-seps in pt-BR and pt-PT almost halved when the report goals are increased. We use noise multiplier 7 for TreeAgg, which is determined by StackOverflow simulation experiments [35]. The noise multiplier of BandMF is computed so that the RMSE is similar to TreeAgg with noise multiplier 7 [8] and will achieve stronger DP guarantees. As BLT achieves comparable privacy-utility trade-off as BandMF in the high single-to-noise ratio regime ($\epsilon=8$) in simulation experiment (Tab. 3 ), the noise multiplier of BLT is computed so that BLT achieves similar DP guarantee as BandMF for the desired round. (2) BLT achieves better privacy-utility trade-off compared to TreeAgg and BandMF.BLTachieves comparable and slightly better NWP accuracy for the models in Tab. 4 and Fig. 7, and stronger DP guarantees. The model performance are further verified by A/B test in the production application comparing BLT models to baseline models for WMR and WPM. The advantage of BLT in privacy-utility trade-off is clearly demonstrated in Fig. 15, and BLT is better than not only TreeAgg, but also BandMF across the production LM training. The practical min-sep can be quite different from the estimated min-seps for optimizing BandMF and BLT matrices, $\sim$90 compared to 100 for pt-PT, $\sim$300 compared to 400 for es-ES, $\sim$440 compared to 400 for id-ID, and 2000+ compared to 1000 for pt-BR. As BLT is more flexible on min-sep estimation, the challenge of reliably estimating min-sep resulting in BLT achieving even stronger privacy-utility trade-offs than BandMF in the production LMs training.
@@ 6 Concluding Remarks
The empirical results in this paper primarily focus on the cross-device FL setting where privacy amplification by sampling is challenging in practice. The discussions (e.g. Tab. 1) can also be applied to centralized setting for user-level DP or example-level DP. In centralized setting, BandMF [8] with amplification can achieve better privacy-utility trade-off measured by RmsLoss among the mentioned mechanisms, when number of rounds $n$ and model dimension $m$ is not too large for optimizing and applying the mechanism. When $n$ and $m$ are large, BLT and BandToep [25] (similarly, BandFHU [24]) can both be applied, where BLT has less optimization cost for very large $n$ (shown in Fig. 1), while BandToep can apply existing amplification by sampling.
@@ Acknowledgement
The authors thank Zachary Garrett, Nicole Mitchell, and Keith Rush for supporting the production code; Ryan McKenna and Yuanbo Zhang for reviewing an early draft.

### v2

@@ Abstract
The state-of-the-art for training on-device language models for mobile keyboard applications combines federated learning (FL) with differential privacy (DP) via the DP-Follow-the-Regularized-Leader (DP-FTRL) algorithm. Two variants of DP-FTRL are used in practice, tree aggregation and matrix factorization. However, tree aggregation suffers from significantly suboptimal privacy/utility tradeoffs, while matrix mechanisms require expensive optimization parameterized by hard-to-estimate-in-advance constants, and high runtime memory costs. This paper extends the recently introduced Buffered Linear Toeplitz (BLT) mechanism to multi-participation scenarios. Our BLT-DP-FTRL maintains the ease-of-use advantages of tree aggregation, while essentially matching matrix factorization in terms of utility and privacy. We evaluate BLT-DP-FTRL on the StackOverflow dataset, serving as a re-producible simulation benchmark, and across four on-device language model tasks in a production FL system. Our empirical results highlight the advantages of the BLT mechanism and elevate the practicality and effectiveness of DP in real-world scenarios.
@@ 2.3 Matrix Factorization (MF) for DP-FTRL
$\widetilde{{\bm{A}}{\bm{X}}}={\bm{B}}({\bm{C}}{\bm{X}}+\zeta{\bm{Z}})={\bm{A}}({\bm{X}}+\zeta{\bm{C}}^{-1}{\bm{Z}})\Rightarrow\tilde{\Delta}^{t}=\Delta^{t}+(\zeta{\bm{C}}^{-1}{\bm{Z}})_{t,:}$ (1)
@@ 2.4 BLT Mechanisms in DP-FTRL
Applying the parameterization in Eq. 2 (also appear in [15, Sec 5]) to [15, Alg 1], and initializing buffers ${\bm{S}}_{-1}\leftarrow\mathbf{0}\in\mathbb{R}^{d\times m}$, we can efficiently compute ${\bm{Z}}_{t,:}$ from $\hat{\bm{Z}}_{t,:}$ in the streaming setting, ${\bm{Z}}_{t,:}=\hat{\bm{Z}}_{t,:}+\omega^{T}{\bm{S}}_{t-1},{\bm{S}}_{t}=\operatorname{diag}(\theta){\bm{S}}_{t-1}+\bm{1}_{d}\hat{\bm{Z}}_{t,:}$. We rearrange the update equations to get $\hat{\bm{Z}}$ from ${\bm{Z}}$ and ${\bm{S}}$,
@@
For completeness, we provide the streaming multiplication algorithm of ${\bm{Z}}={\bm{C}}\hat{\bm{Z}}$ and $\hat{\bm{Z}}=C^{-1}{\bm{Z}}$ for ${\bm{C}}=\text{BLT}(\theta,\omega)$ in Alg. 2 and Alg. 3, respectively. Alg. 2 is a direct application of [15, Alg 1] and only used to derive Alg. 3, which is our streaming algorithm for generating correlated noise with BLTs using only $dm$ memory. Finally, we apply the streaming correlated noise $\hat{\bm{Z}}$ by BLT from Alg. 3 (using Eq. 3) in Alg. 1 for the BLT-DP-FTRL algorithm, i.e.,
@@
       $\theta\in\mathbb{R}^{d},\omega\in\mathbb{R}^{d}$ for ${\bm{C}}=\text{BLT}(\theta,\omega)$
@@
  Initialize buffers ${\bm{S}}_{-1}\leftarrow\mathbf{0}\in\mathbb{R}^{d\times m}$
@@
       $\theta\in\mathbb{R}^{d},\omega\in\mathbb{R}^{d}$ for ${\bm{C}}=\text{BLT}(\theta,\omega)$
@@
  Initialize buffers ${\bm{S}}_{-1}\leftarrow\mathbf{0}\in\mathbb{R}^{d\times m}$
@@ 3.1 Sensitivity Under Multiple Participations
The sensitivity of general Toeplitz matrices with decaying coefficients is recently discussed in Kalinin and Lampert [26, Thm. 2], which we restate in Thm. 3.1 with our notation. The participation pattern $\pi^{\star}$ simply puts the $k$ participations as early as possible, with each participation separated by exactly $b$. This sensitivity computation is important for both DP accounting and optimizing for BLT parameters in Sec. 3.3.
@@ 3.2 Evaluating Matrix Factorization Mechanisms
Eqs. 9 and 10 measure the distribution of noise to approximate the privacy-utility trade-off as the loss, which do not depend on the noise multiplier $\alpha=\sigma/\operatorname{sens}_{\Pi}({\bm{C}})$ that is directly used in accounting for the DP guarantees, as discussed in [15, Introduction]. For optimized ${\bm{C}}$ in matrix factorization mechanisms (e.g., TreeAgg, BandMF, and BLT), specific DP guarantees are achieved by scaling $\alpha$ (and corresponding $\sigma$ in Alg. 1). The total noise on the prefix sum $BZ$ also scales MaxLoss and RmsLoss by $\alpha$. Hence, without loss of generality, we use RmsLoss and MaxLoss to compare the optimality of different matrix factorization mechanisms, which is equivalent to assuming $\alpha=1$ that corresponds to $(\epsilon=5.3,\delta=10^{-7})$-DP (for example).
Note we deviate from the definitions of [15, arXiv v3], in order to distinguish the error introduced by ${\bm{B}}$ from the total loss (which is what we optimize and use to compare mechanisms), which incorporates the sensitivity of ${\bm{C}}$.
@@ 3.3 Optimizing BLTs for Multiple Participations
  $\triangleright$ Note that in [15], ${\bm{C}}$ was parameterized by $\hat{\theta}$ and ${\bm{C}}^{-1}$ was parameterized by $\theta$.
  $p(x)=\prod_{i\in[d]}(1-\hat{\theta}_{i}x)$
  $q(x)=\prod_{i\in[d]}(1-\theta_{i}x)$
@@
For the typical scale of $n<10^{5}$ in FL systems, rather than deriving closed forms for sensitivity and error as in Dvijotham et al. [15], we use an alternative approach that is flexible and simple to implement. Recall the properties of BLTs discussed in Sec. 2.4, we parameterize the optimization of the BLT by the pair $(\theta,\hat{\theta})$. Dvijotham et al. [15, Lem. 5.2] implies given a pair $(\theta,\hat{\theta})$, there exist unique $(\omega,\hat{\omega})$ such the $\text{BLT}(\theta,\omega)^{-1}=\text{BLT}(\hat{\theta},\hat{\omega})$, and we can compute $\omega$ and $\hat{\omega}$ in time $\mathcal{O}(d^{2})$; this result is summarized below as Alg. 5. Thus, given a $(\hat{\theta},\theta)$, we can efficiently compute the Toeplitz coefficients of ${\bm{C}}$ (using Eq. 2 applied to $(\theta,\omega)$) and ${\bm{C}}^{-1}$ (applying Eq. 2 to $(\hat{\theta},\hat{\omega})$). From the Toeplitz coefficients of ${\bm{C}}$ we can then efficiently compute sensitivity using Thm. 3.1. RmsError can be computed efficiently from the Toepltiz coefficients of ${\bm{C}}^{-1}$ following the approach of McKenna [27, Prop. 3.1], and a simple generalization of this approach applies to to MaxError as well. For completeness we summarize in the following proposition:
@@ 5 Production LMs for Mobile Keyboard
We summarize the privacy and utility results for es-ES, id-ID, pt-BR, and pt-PT LMs in Tab. 4, and show the privacy-utility curves for training pt-PT in Fig. 7. We provide additional curves for other LMs in Figs. 13, 14 and 15 inApp. B, and discuss the observations here. (1) Most of the models achieve DP guarantee of $\epsilon<10$ with exception of $\epsilon\sim 10$ for pt-PT due to the challenge of small population; the pt-BR model trained with BLT-16.1 achieves $\epsilon<1$ at round 2000.DP guarantees of $\epsilon<10$ is commonly used for machine learning, and $\epsilon<1$ are considered strong guarantees [31]. To achieve single-digit DP guarantees in practice without sacrificing the utility, the production LMs are trained with large number of clients per round (known as report goal in FL systems). We use typical report goal 6.5K [38] for es-ES, id-ID and pt-BR, and use a smaller report goal 3K for pt-PT with a smaller population. We additionally run BLT-DP-FTRL with larger report goal and linearly scale up the noise multiplier to keep the signal-to-noise ratio for utility: BLT-16.1 with report goal 12K for pt-BR and BLT-5.5 with report goal 6K for pt-PT. The resulting min-seps in pt-BR and pt-PT almost halved when the report goals are increased. We use noise multiplier 7 for TreeAgg, which is determined by StackOverflow simulation experiments [38]. The noise multiplier of BandMF is computed so that the RMSE is similar to TreeAgg with noise multiplier 7 [8] and will achieve stronger DP guarantees. As BLT achieves comparable privacy-utility trade-off as BandMF in the high single-to-noise ratio regime ($\epsilon=8$) in simulation experiment (Tab. 3 ), the noise multiplier of BLT is computed so that BLT achieves similar DP guarantee as BandMF for the desired round. (2) BLT achieves better privacy-utility trade-off compared to TreeAgg and BandMF.BLTachieves comparable and slightly better NWP accuracy for the models in Tab. 4 and Fig. 7, and stronger DP guarantees. The model performance are further verified by A/B test in the production application comparing BLT models to baseline models for WMR and WPM, where we target on improved or neutral utilities. The advantage of BLT in privacy-utility trade-off is clearly demonstrated in Fig. 15, and BLT is better than not only TreeAgg, but also BandMF across the production LM training. The practical min-sep can be quite different from the estimated min-seps for optimizing BandMF and BLT matrices, $\sim$90 compared to 100 for pt-PT, $\sim$300 compared to 400 for es-ES, $\sim$440 compared to 400 for id-ID, and 2000+ compared to 1000 for pt-BR. As BLT is more flexible on min-sep estimation, the challenge of reliably estimating min-sep resulting in BLT achieving even stronger privacy-utility trade-offs than BandMF in the production LMs training.
@@
Model launches
We have launched several language models for on-the-fly (OTF) rescoring words in Gboard smart completion and suggestion [36, 38]. The models trained with BLT-DP-FTRL algorithm in this paper achieved better privacy-utility trade-off than the TreeAgg-DP-FTRL algorithm [25] widely used in FL practice [38], and comparable with MF-DP-FTRL [8] that achieves state-of-the-art performance and some $\epsilon<1$ models reported in [36]. We report $\rho$-zCDP and $(\epsilon,\delta)$-DP of $\delta=10^{-10}$ following Xu et al. [38], Xu and Zhang [36]; the details of the DP configuration are provided in App. C. The Portuguese OTF LM in Brazil (pt-BR) is launched with DP $\epsilon=1.234$ (alternatively, zCDP $\rho=0.022$); pt-PT OTF LM is launched with DP $\epsilon=8.19$ (alternatively, zCDP $\rho=0.749$) compared to DP $\epsilon=9.56$ (alternatively, zCDP $\rho=0.99$) in [38], and significant improvement in WMR of $-0.91\%$; es-ES OTF LM is launched with DP $\epsilon=4$ (alternatively, zCDP $\rho=0.202$) compared to DP $\epsilon=9.01$ (alternatively, zCDP $\rho=0.89$) in [38]. In addition, the discussed default BLT mechanism optimized for $b=400$ and noise multiplier 7.379 is used to train and launch more models: the English LM is launched with DP $\epsilon=4.84$ (alternatively, zCDP $\rho=0.29$); the Spanish LM in Argentina is launched with DP $\epsilon=5.73$ (alternatively, zCDP $\rho=0.392$); the Spanish LM in Mexico is launched with DP $\epsilon=11.58$ (alternatively, zCDP $\rho=1.39$). The id-ID model is actively trained and tested with new vocabulary constructed from [32]. The Gboard OTF LMs are trained with FL and DP following the commitment in [36, 40], and we made significant progress on upgrading all existing FL LMs with advanced techniques for DP.
@@ 6 Concluding Remarks
The empirical results in this paper primarily focus on the cross-device FL setting where privacy amplification by sampling is challenging in practice. The discussions (e.g. Tab. 1) can also be applied to centralized setting for user-level DP or example-level DP. In centralized setting, BandMF [8] with amplification can achieve better privacy-utility trade-off measured by RmsLoss among the mentioned mechanisms, when number of rounds $n$ and model dimension $m$ is not too large for optimizing and applying the mechanism. When $n$ and $m$ are large, BLT and BandToep [27] (similarly, BandFHU [26]) can both be applied, where BLT has less optimization cost for very large $n$ (shown in Fig. 1), while BandToep can apply the known amplification by sampling mechanism. In conclusion, BLT-DP-FTRL algorithm is recommended for cross-device federated learning [38, 24], and is also useful for DP in datacenter [31, 37, 7] and federated learning in trusted execution environments (TEEs) [11, 19].
@@ Acknowledgement
The authors thank Zachary Garrett, Nicole Mitchell, and Keith Rush for supporting the production code; Yu Xiao for helping some of the production experiments; Ryan McKenna and Yuanbo Zhang for reviewing an early draft; and Nikita Kalinin and Krishna Pillutla for pointing out typos in Alg. 2 and Alg. 5 in the initial arXiv version.
@@ Appendix B Additional Simulation and Production Results
Appendix C Reporting privacy guarantees
This section clarifies the nuances of the DP guarantees following the guidelines outlined in [31, Sec. 5.3]. They are similar (if not identical) to [38, Appendix A] and [35, Appendix C.3], except for the exact DP guarantees. We include the entire checklist for completeness.
  1. 1.
    DP setting. This is a central DP guarantee where the service provider is trusted to correctly implement the mechanism.
  2. 2.
    Instantiating the DP Definition
    1. (a)
      Data accesses covered: The DP guarantee applies to all well-behaved clients [^3] Clients that faithfully follow the algorithm including participation limits. Due to the design of the algorithm, a mis-behaved client does not adversely affect the DP guarantee of any well-behaved clients. in a single training run. We do not account for hyperparameter tuning, or the selection of the final model checkpoint using evaluation metrics or A/B testing in our guarantees. Public data such as C4 [39] or LLM-based synthetic data [35], are used for pre-training.
    2. (b)
      Final mechanism output: Only the final model checkpoint is released for production deployment, however the mechanism’s output is technically the full sequence of privatized gradients, and so the guarantee also applies at this level, and hence all intermediate models are protected (including those sent to devices participating in federated learning).
    3. (c)
      Unit of privacy. Device-level DP is considered, i.e., the notion of adjacency is with respect to arbitrary training datasets on each client device, and the device might have an arbitrarily large local dataset containing arbitrary training examples. For user’s with a single device, this corresponds directly to user-level DP; for devices shared with multiple users, this provides a stronger notion of DP than user-level; for a user with multiple devices that happen to both participate in training the model, the notion is weaker, but group privacy can be used to obtain a user-level guarantee.
    4. (d)
      Adjacency definition for “neigbouring” datasets: We use the zero-out definition [25]. This is a a special form of the add-or-remove definition, where neighboring data sets differ by addition/removal of a single client. In the absence of a client at any training step, we assume that the client’s model update gets replaced with the all zeros vector. This assumption enforces a subtle modification to the traditional definition of the add/remove notion of DP which allows neighboring data sets to have the same number of records.
  3. 3.
    Privacy accounting details
    1. (a)
      Type of accounting used: Both $\rho-$zCDP [5] accounting, and PLD accounting [14] for $(\epsilon,\delta)-$DP are used.
    2. (b)
      Accounting assumptions : Each client only participates limited times during the training, and there are at least a min-separation number of rounds between two consecutive participation of a client. Client participation is enforced by a timer on clients in the cross-device FL system.
    3. (c)
      The formal DP statement: See Sec. 5 model launches for reported DP guarantees.
    4. (d)
      Transparency and verifiability: We use the open-sourced core implementation code in TensorFlow Federated and Tensorflow Privacy. Key portions of the cross-device FL system are also open sourced.